OpenStudy (anonymous):

using base 16 multiply 1C x 3 in the given base

OpenStudy (anonymous):

I would do it in base 10 and then convert the answer back to base 16 In base 16 C is 12 (A is 10, B is 11, C is 12, D is 13, etc...) 1C = 1x16^1 + 12x16^0 = 16 + 12 = 28 ... in base 10 3 x 28 = 84 Now convert 84 to base 16 84 = 5*16^1 + 4*16^0 = 54 ... in base 16

OpenStudy (anonymous):

Thank you! This is supposed to be a class that utilizes everyday math... not sure when ill ever use this!